OpenStudy (anonymous):

solve using the addition and multipication principles. x/6 - 1 < less than or equal to 1/3

OpenStudy (anonymous):

can some help me with this problem please solve using the addition and multiplication principles. x/6 - 1 < less than or equal to 1/3

OpenStudy (gw2011):

(x/6) - 1 <= 1/3 The symbol <= means less than or equal to (this is the only way that I could write it--the proper way of writing it is < with one horizontal line under <) x - 6 <= 2 Bothe sides were multiplied by 6 to eliminate the denominator 6 x<= 8